the sum of two numbers is -9 and their difference is 31, what are the numbers?
Answer by addingup(3677)

x+y = -9 (1)
x-y = 31 here, add y on both sides
x = 31+y using this value for x, substitute in (1)
31+y+y = -9
2y = -40
y = -20
and since x-y = 31:
x-(-20) = 31
x+20 = 31
x = 11
--------------------
Check:
x+y = -9
11+(-20) = -9 correct
x-y = 31
11-(-20) = 11+20 = 31 also correct